OpenStudy (anonymous):

who wanna help me plot these x+yor equal to 0 y>or equal to 0

OpenStudy (welshfella):

for the first one plot the graph of x + y = 12 can you do that?

OpenStudy (welshfella):

it is easy note that when x = 0 y =12 and when y = 0 x = 12 so plot the point (0,12) and (12,0) and join them up It well be a straight line.

OpenStudy (welshfella):

the for x + y <= 12 shade the area below the line

OpenStudy (anonymous):

sorry i still dont understand its suppose to \[y=\le2x+3\]

OpenStudy (anonymous):

but is the plot 0,12

OpenStudy (welshfella):

i described is the plot of x + y <= 12

OpenStudy (anonymous):

no \[y=\le2x+3\]

OpenStudy (welshfella):

plot the line 2x + 3 first plot 3 points - they should be in a straight line if youhave done them correctly one point would be when x = 0:- y = 2(0) + 3 = 3 so this point will be (0,3) plot 2 more

OpenStudy (anonymous):

what about this one \[y \le2x+3\]

OpenStudy (welshfella):

this one will be the areas below the the line y = 2x + 3 - shade it in.

OpenStudy (welshfella):

it will look something like this:- 

OpenStudy (anonymous):

lol ok but what's the plot it's hard for me to make plots without numbres on the graph

OpenStudy (welshfella):

well you cant plot points accurately without numbers on the graph. Haven't you got graph paper?

OpenStudy (anonymous):